OpenStudy (anonymous):

If tan(x) = -1/3, cos(x) >0, then what is sin(2x), cos(2x), and tan(2x)?

OpenStudy (aum):



OpenStudy (aum):

x must be in the fourth quadrant where tan(x) is negative and cos(x) is positive. Find the hypotenuse in the diagram shown above. Then find sin(x), cos(x). sin(x) should be negative and cos(x) should be positive. sin(2x) = 2sin(x)cos(x). Substitute and find sin(2x).

OpenStudy (aum):

cos(2x) = 2cos^2(x) - 1 tan(2x) = sin(2x) / cos(2x)

OpenStudy (xapproachesinfinity):

 this is your triangle in the fourth quadrant

OpenStudy (xapproachesinfinity):

like @aum said find the hypotenuse first then find sinx conx tanx and use them for double angles

OpenStudy (anonymous):

and so the hypotenuse would be sqrt(3^2+1^2) , which gives about 3.16

OpenStudy (xapproachesinfinity):

yes! but you don't need decimal \(\sqrt{10}\) leave it this way

OpenStudy (anonymous):

so sin(x) = -1/sqrt(10) , cos(x) = 3/sqrt(10) , tan(x) = -1/3. right?

OpenStudy (aum):

Correct.

OpenStudy (aum):

sin(2x) = 2sin(x)cos(x) = 2 * (-1 / sqrt(10) ) * (3 / sqrt(10)) = ?

OpenStudy (anonymous):

that would be -3/5

OpenStudy (aum):

Yes.

OpenStudy (anonymous):

cos(2x) would be 4/5

OpenStudy (aum):

correct.

OpenStudy (anonymous):

and tan(2x) would be -3/4

OpenStudy (aum):

Got it.

OpenStudy (anonymous):

Thank you so much!

OpenStudy (aum):

You are welcome.